OpenStudy (anonymous):

Help me with some questions? Given f(x) = 6x2 – x – 12 and g(x) = 2x – 3, find the function (fg)(x). Evaluate the composite function f(g(x)) for x = 31.

OpenStudy (anonymous):

Much gratitude to whoever helps me. I just can't get these. I hate alg 2

terenzreignz (terenzreignz):

f + g ? You just add f(x) and g(x)

OpenStudy (anonymous):

would the answer just be x^2? for that one?

terenzreignz (terenzreignz):

As a matter of fact, yes :D

terenzreignz (terenzreignz):

now the composite function f(g(x)) is a little tricky What you do is replace all instances of x in f(x) with g(x)

terenzreignz (terenzreignz):

\[\large f(\color{orange}x) =2 - \color{orange}x^2\]\[\large \color{red}{g(x)=3x+10}\] \[\Large f[\color{red}{g(x)}]=2-[\color{red}{g(x)}]^2\]

OpenStudy (anonymous):

And then I just solve that?^

terenzreignz (terenzreignz):

replace g(x) and then solve from there, yes :D

OpenStudy (anonymous):

okay I get x-5. Is that correct?

terenzreignz (terenzreignz):

I don't think so. Remember \[\large g(x) = 3x+10\]

OpenStudy (anonymous):

hmmm, is it 6x+20?

terenzreignz (terenzreignz):

You're guessing. Replace the g(x) here \[\Large f[\color{red}{g(x)}]=2-[\color{red}{g(x)}]^2\] with 3x + 10 like so... \[\Large f[\color{red}{g(x)}]=2-[\color{red}{3x+10}]^2\]

terenzreignz (terenzreignz):

and carry on from there...

OpenStudy (anonymous):

No i'm not. Do I just combine? Coz what I did was multiplied within the parentheses.

terenzreignz (terenzreignz):

First, you evaluate \[\Large [3x+10]^2\]

OpenStudy (anonymous):

6x+20 right?

OpenStudy (anonymous):

9x sorry

OpenStudy (anonymous):

ugh wait 9x+100

OpenStudy (anonymous):

@terenzreignz

terenzreignz (terenzreignz):

Could you review the FOIL method or the method of squaring a binomial? Or, as a quick reference, look at this... \[\Large (\color{red}a+\color{blue}b)^2 = \color{red}a^2 + 2\color{red}a\color{blue}b+\color{blue}b^2\]

terenzreignz (terenzreignz):

By that logic, what is \[\Large (\color{red}{3x}+ \color{blue}{10})^2= \qquad?\]

OpenStudy (anonymous):

ahhh okay 3^2+2(3)(10)+10^2 like that?

terenzreignz (terenzreignz):

Do remember the x... your 'a' in this case is not simply 3 but 3x

OpenStudy (anonymous):

okay 3x^2+2(3x)(10)+10^2

OpenStudy (anonymous):

then combine?

terenzreignz (terenzreignz):

\[\large3x^2\qquad \color{red}?\]

OpenStudy (anonymous):

I don't understand what you're asking

terenzreignz (terenzreignz):

Why is it 3x^2 ? What is the square of 3x ?

OpenStudy (anonymous):

oh 9x^2

terenzreignz (terenzreignz):

better. Okay, so simplify?

OpenStudy (anonymous):

okay 3x^2+2(3x)(10)+10^2= 9x+6x+5+100

OpenStudy (anonymous):

or I mean 105?

terenzreignz (terenzreignz):

Why do you have 3x^2 again? -.-

OpenStudy (anonymous):

no I changed it to 9x+6x+105

terenzreignz (terenzreignz):

let's start again... \[\Large (3x)^2+2(3x)(10)+10^2\] And simplify this one at a time, please.

OpenStudy (anonymous):

okay so (3x)^2=9x

terenzreignz (terenzreignz):

9x^2 !!!

OpenStudy (anonymous):

where are you getting 9x^2 from? but anyways 81x

terenzreignz (terenzreignz):

No... I mean, you square both the 3 and the x -.- \[\Large (3x)^2=\color{red}{9x^2}\]

OpenStudy (anonymous):

oh okay. Got it

OpenStudy (anonymous):

9x^2+6x+5+100x^2 better?

terenzreignz (terenzreignz):

Nope. Now what about this part (in red)? It's just multiplication. \[\Large9x^2 + \color{red}{2(3x)(10)}+10^2\]

OpenStudy (anonymous):

sorry 6x+20 right?

terenzreignz (terenzreignz):

wrong. what's 2(3x)(10)

OpenStudy (anonymous):

60x

terenzreignz (terenzreignz):

right. now this. \[\Large 9x^2 + 60x +\color{red}{10^2}\]

OpenStudy (anonymous):

100

terenzreignz (terenzreignz):

\[\Large 9x^2 +60x + 100\] And this is what we put in this place... \[\Large f[g(x)]= 2 - \color{red}{[g(x)]^2}\]

OpenStudy (anonymous):

omg sorry for taking up so much of your time

terenzreignz (terenzreignz):

you won't be for longer, lol, I have to get to class in a few minutes

OpenStudy (anonymous):

ah okay thanks anyways!

terenzreignz (terenzreignz):

I'm sure there are other people on OS who can help you out, but unfortunately, for the moment, I have to go :D -------------------------------------------- Terence out

OpenStudy (anonymous):

Still need help with this problem?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

the person answered a different question before this

OpenStudy (anonymous):

this problem? Given f(x) = 6x2 – x – 12 and g(x) = 2x – 3, find the function (fg)(x). Evaluate the composite function f(g(x)) for x = 31.

OpenStudy (anonymous):

before we start though, do you know what g(2) = ?

OpenStudy (anonymous):

no they didn't answer that. and no..

OpenStudy (anonymous):

g(2) is what you get when you plug 2 for x in g(x)

OpenStudy (anonymous):

g(2) = 2(2) -3 = 1

OpenStudy (anonymous):

since g(x) = 2x - 3

OpenStudy (anonymous):

so f(g(x)) is the same principle. we plug in g(x) in every x in f(x), same way that g(2) is plugging 2 in every x of g(x)

OpenStudy (anonymous):

so f(g(x)) = 6[g(x)]^2 - g(x) - 12 do you follow so far?

OpenStudy (anonymous):

Ohh okay

OpenStudy (anonymous):

so we have: \[f(g(x)) = 6(2x - 3)^2 - (2x -3) -12\] knowing the shortcut, by habit, that: \[(a+b)^2 = a^2 + 2ab + b^2\] i'll save my self a small step and a small mess to expand. \[f(g(x)) = 6(4x^2 - 12x + 9) - 2x + 3 - 12 = 24x^2 -74x + 45\] that's (fg)(x). now they ways (fg)(31) which is equivalent to f(g(31)) [diffrent notation] plug in 31 for every x in f(g(x))

OpenStudy (anonymous):

now they want**

OpenStudy (anonymous):

:/ oof this..problem looks intimidating

OpenStudy (anonymous):

it's easier than it seems. and with a bit of practice they will become very easy and second nature. i have may have complicated things with the "shortcut". expanding with foil is as good. don't let math intimidate you, it's your friend :) and we're here to help you love it!

OpenStudy (anonymous):

okay so with every x in there I replace is with a 31? @Euler271

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

okay 6(4(31)^2-12(31)+9)-2(31)+3-12=24(31)^2+74(31)+45

OpenStudy (anonymous):

@Euler271

OpenStudy (anonymous):

ya they would be the same thing

OpenStudy (anonymous):

Um for the first part I get 20815... @Euler271

OpenStudy (anonymous):

i get 25403

OpenStudy (anonymous):

oh okay the 2nd part I get 25403 too

OpenStudy (anonymous):

@Euler271

OpenStudy (anonymous):

cool ^_^ thats the final answer

OpenStudy (anonymous):

Oh! @euler. those were both two different problems though...

OpenStudy (anonymous):

@euler this was for the 2nd problem